# And Then the Hammer Broke: Seeing Machine Vision

### 09 Mar 2024 v1

@@ 1 The Hammer
Looking at those colorful photos of vast outer planets hundreds of millions of miles away captured by giant telescopes, one might begin to believe that really are seeing the true majesty of the celestial bodies. These images, Haraway writes, appear to “let the viewer ‘experience’ the moment of discovery in immediate vision of the ‘object’.”[5, 582] Likewise, imaging devices capture the microscopic world at wavelengths imperceivable to humans, opening up a mysterious microbiological world of “defending T cells and invading viruses.”[5, 582] Wielding the Heideggerian hammer, we cease to become aware of the instruments of vision and thereby buy into the god trick: both that we can see everything in the universe as they really are, and that this view is from ‘nowhere’, in that my own vision may be subjectively embodied but my instruments of vision allow me to access objective vision ‘beyond’ my body. Importantly, the instruments of vision are not the ‘source’ or ‘constitutor’ of the god trick. A digital telescope is ‘merely’ a technological object which records radiation from space, and microscopic imaging techniques ‘merely’ ping organic structures with radiation to collect data about their density and organization. That is to say: they are material machines. It is us humans, like the carpenter, who use the outputs of these machines to generate fantasies of knowledge – admittedly beautiful fantasies – of the cosmos’ grand beauty, and of the countless miniature wars fought within the very matter of our body. This is not only observed in the scientific domain, but in our world of media. Scrolling through infinite video feeds on the Internet, one may subscribe to the illusion that they are seeing the world as it is, ceasing consciousness of the army of cameras – pointed in particular directions by particular people – which produced each of those videos. Haraway understands that these instruments are Heideggerian hammers, that they break: “The only people who end up actually believing and, goddess forbid, acting on the ideological doctrines of disembodied scientific objectivity… are nonscientists.”[5, 576] The astronomer who runs the computer software which converts raw radiation signals into “coffee table color photographs”[5, 583] is not under the illusion that we are looking at the cosmos when we peer into those images, nor that the fantasies which spring from those images are imposable back onto the ‘real’ universe. Haraway is right, then, when she notes that scientists “might be our allies.”[5, 594]
@@
But I worry that the examples raised in this 1988 text of Haraway’s, “Situated Knowledges”, may not itself be sufficiently situated in our present (scientific and technological) knowledge. Machine vision, I argue, does not participate in the god trick in the same way anymore. Haraway’s machines – the microscopes and telescopes – are content with being instruments wielded by humans. But contemporary machine vision aspires for much more. We can get a sense for this shift by understanding how the researchers that built our contemporary technologies of machine vision understand (machine) vision. Li Fei-Fei and Ranjay Krishna, two prominent researchers in the field, take vision to be “the ability to see” and as “central to intelligence”.[3, 86] Central to their analysis is that “computer vision, like human vision, is not just perception; it is deeply cognitive.”[3, 95] The first monumental goal for computer vision as a budding research area in the late 20th century was the recognition of particular objects in a visual field, such as a house or a dog. This is a trivial task for most humans, but for a computer, an image is merely a collection of numbers arranged along height, width, and color axes: so, from an algorithmic perspective, there is no immediately clear association of pixels which indicates the presence of a house, a dog, or some other category. To adapt, Fei-Fei and Krishna write, machine vision research shifted its focus from algorithms to data. esearchers had initially hand-designed algorithms to check for particular patterns and important features in the image, but these failed to capture the variability and dynamic nature of object categories. Machine vision started the “search for a new approach with one key assumption: even the best algorithm would not generalize well if the data it learned from did not reflect the real world.”[3, 87] This is a significant paradigm shift: building machine vision is no longer principally about building a better machine, but rather about building a better map of the world. This map, Fei-Fei and Krishna suggest, must achieve three design goals: “scale (a large quantity of data), diversity (a rich variety of objects), and quality (accurately labeled objects).”[3, 88] The instances of machine vision that we encounter in our current world – the New York facial recognition model, Chinese psychological profiling systems, autonomous war drones, but also social media filters and virtual meeting software – are farther from Haraway’s telescopes and closer to this new kind of machine vision.

### 08 Apr 2024 v2

@@ 1 The Hammer
Looking at those colorful photos of vast outer planets hundreds of millions of miles away captured by giant telescopes, one might begin to believe that really are seeing the true majesty of the celestial bodies. These images, Haraway writes, appear to “let the viewer ‘experience’ the moment of discovery in immediate vision of the ‘object’.”[5, 582] Likewise, imaging devices capture the microscopic world at wavelengths imperceivable to humans, opening up a mysterious microbiological world of “defending T cells and invading viruses.”[5, 582] Wielding Thor’s hammer, we cease to become aware of the instruments of vision and thereby buy into the god trick: both that we can see everything in the universe as they really are, and that this view is from ‘nowhere’, in that my own vision may be subjectively embodied but my instruments of vision allow me to access objective vision ‘beyond’ my body. Importantly, the instruments of vision are not the ‘source’ or ‘constitutor’ of the god trick. A digital telescope is ‘merely’ a technological object which records radiation from space, and microscopic imaging techniques ‘merely’ ping organic structures with radiation to collect data about their density and organization. That is to say: they are material machines. It is us humans, like the carpenter, who use the outputs of these machines to generate fantasies of knowledge – admittedly beautiful fantasies – of the cosmos’ grand beauty, and of the countless miniature wars fought within the very matter of our body. This is not only observed in the scientific domain, but in our world of media. Scrolling through infinite video feeds on the Internet, one may subscribe to the illusion that they are seeing the world as it is, ceasing consciousness of the army of cameras – pointed in particular directions by particular people – which produced each of those videos. Haraway understands that these instruments are Heideggerian hammers, that they break: “The only people who end up actually believing and, goddess forbid, acting on the ideological doctrines of disembodied scientific objectivity… are nonscientists.”[5, 576] The astronomer who runs the computer software which converts raw radiation signals into “coffee table color photographs”[5, 583] is not under the illusion that we are looking at the cosmos when we peer into those images, nor that the fantasies which spring from those images are imposable back onto the ‘real’ universe. Haraway is right, then, when she notes that scientists “might be our allies.”[5, 594]
@@
But I worry that the examples raised in this 1988 text of Haraway’s, “Situated Knowledges”, may not itself be sufficiently situated in our present (scientific and technological) knowledge. Machine vision, I argue, does not participate in the god trick in the same way anymore. Haraway’s machines – the microscopes and telescopes – are content with being instruments wielded by humans. But contemporary machine vision aspires for much more. We can get a sense for this shift by understanding how the researchers that built our contemporary technologies of machine vision understand (machine) vision. Li Fei-Fei and Ranjay Krishna, two prominent researchers in the field, take vision to be “the ability to see” and as “central to intelligence”.[3, 86] Central to their analysis is that “computer vision, like human vision, is not just perception; it is deeply cognitive.”[3, 95] The first monumental goal for computer vision as a budding research area in the late 20th century was the recognition of particular objects in a visual field, such as a house or a dog. This is a trivial task for most humans, but for a computer, an image is merely a collection of numbers arranged along height, width, and color axes: so, from an algorithmic perspective, there is no immediately clear association of pixels which indicates the presence of a house, a dog, or some other category. To adapt, Fei-Fei and Krishna write, machine vision research shifted its focus from algorithms to data. Researchers had initially hand-designed algorithms to check for particular patterns and important features in the image, but these failed to capture the variability and dynamic nature of object categories. Machine vision started the “search for a new approach with one key assumption: even the best algorithm would not generalize well if the data it learned from did not reflect the real world.”[3, 87] This is a significant paradigm shift: building machine vision is no longer principally about building a better machine, but rather about building a better map of the world. This map, Fei-Fei and Krishna suggest, must achieve three design goals: “scale (a large quantity of data), diversity (a rich variety of objects), and quality (accurately labeled objects).”[3, 88] The instances of machine vision that we encounter in our current world – the New York facial recognition model, Chinese psychological profiling systems, autonomous war drones, but also social media filters and virtual meeting software – are farther from Haraway’s telescopes and closer to this new kind of machine vision.